### v3

@@ 1. Introduction
Street intersections—the junctions where two or more streets segments meet—are important transport network features whose configuration contributes to urban connectivity, centrality, compactness, walkability, and safety (Fogliaroni et al.,, 2018; Sharifi,, 2019; Cerin et al.,, 2022). Despite street intersections’ apparent simplicity, counting them can be nuanced and challenging in practice (Yang et al.,, 2020; Li et al.,, 2020). As we will argue, this is primarily due to three modeling challenges: network nonplanarity, intersection complexity, and curve digitization. In turn, intersection miscounts can bias downstream indicators of urban form, but this bias has not been fully accounted, no easily reusable algorithms exist to mitigate it, and intersection counts often remain ad hoc and naïve in research and practice.
This article measures this problem’s extent and asks: can we algorithmically simplify a street network graph’s topology to produce a more accurate and performant model? To answer this, it introduces and validates the OSMnx package’s reusable algorithms that ensure that graph nodes represent individual intersections or dead-ends and that graph edges represent individual street segments. This contribution offers three main advantages over existing techniques: 1) more accurate measures including intersection counts/densities, street segment lengths, node degrees, and betweenness centralities; 2) a topologically-corrected model that better represents the real world; and 3) information compression (i.e., fewer nodes/edges than in the original model) that can drastically speed-up subsequent network algorithms that scale with node or edge count.
@@ 2.1. Street Network Models
then we can say for $G$ that: 1) edge $(u,v)$ links from node $u$ to node $v$, 2) edge $(u,v)$ is incident to $u$ and to $v$, and 3) $v$ is a successor of $u$.
@@ 3.1. Edge Simplification Algorithm
OSMnx’s edge simplification algorithm solves the curve digitization problem, such as occurs with OpenStreetMap raw data as discussed earlier. It simplifies the graph’s topology by removing all nodes that are not “true” edge endpoints (i.e., intersections or dead-ends), then creating new (simplified) edges directly between those true endpoints and retaining the full geometries of the original edges as attributes on those new edges. The full algorithm can be found in the appendix, but we summarize it here.
@@
    If configured, either a) the node possesses a user-defined attribute explicitly denoting it as an endpoint, or b) the node’s incident edges possess a user-defined attribute and have different values than each other in that attribute.
@@ 3.3. Assessment and Validation
To validate the (topological) node consolidation algorithm, for each of these sampled urban areas we randomly sample 120 graph nodes, divided evenly between consolidated and not-consolidated nodes. Each node’s consolidation (or not) result is then validated by qualitatively cross-referencing it, by hand, against Google Street View and satellite imagery data. In comparing these resources, the validator asks whether the node(s) in question would be considered a single intersection by a pedestrian or motorist at that intersection. There are four categories of node consolidation validation result: 1) “true positive,” where the node should be consolidated and is, 2) “true negative,” where the node should not be consolidated and is not, 3) “false positive,” where the node should not be consolidated but is, and 4) “false negative,” where the node should be consolidated but is not. In keeping with the spirit of fuzzy sets, within the “true positive” category we create a subcategory for “true positive but insufficiently robust.” For example, if a sampled validation node is in a roundabout comprising five nodes that should have been consolidated into one, but the algorithm only consolidated four of those five, then this sampled node is assigned this result subcategory. In other words, this result improves on the original model and is mostly correct, but incomplete.
@@ 4. Validation and Assessment Results
Now the question becomes: are these reductions in node count legitimate, or merely false positives being incorrectly consolidated? To that end, the node consolidation algorithm’s validation can be considered from two perspectives, 1) the validity of the algorithm itself, and 2) the validity of its parameterization here. From the first perspective, this deterministic algorithm unsurprisingly performs with 100% accuracy—that is, each sampled node’s result aligns with expectations given this parameterization. From the second perspective, the algorithm performs imperfectly, but better than expected given this simplified constant parameterization per urban area used for global scalability in this preliminary demonstration.
Among nodes that were consolidated (i.e., the “positives”), the validation reveals a 98% true positive rate and a 2% false positive rate. Among these true positives, around 5% were “insufficiently robust,” meaning the algorithm was correct in consolidating a complex intersection but failed to fully merge in all the nodes that the validator deemed integral during their qualitative assessment. Among nodes that were not consolidated (i.e., the “negatives”), the validation reveals an 87% true negative rate and a 13% false negative rate. In other words, our simplified parameterization here was likely overly conservative, leading to more false negatives than false positives through what is essentially algorithmic underfitting. This suggests that the extent of the intersection miscount problem could be even greater than our conservative assessment reveals.
@@ 5. Toward Better Models and Measures
Counting seems simple because its mathematics are simple. But in reality it can be difficult to define what to count and how to count it. Street network intersections exemplify this challenge. In transport planning, innumerable models, measures, and standards—from sustainable development certifications to resilience simulations—rely on intersection counts as input data. Accurate intersection densities, node degrees, and street segment lengths all depend on accurate street network models. Yet modeled intersections can be surprisingly difficult to count accurately, resulting in biased measures that propagate downstream. This study identified the sources of these biases, measured their prevalence, and presented and validated algorithms to better model street intersections for better data-driven, evidence-informed transport planning. These algorithms are generalizable but their implementation here is tied to one data source (OpenStreetMap) or others matching its structure. These open-source, freely resuable methods produce network models with a closer 1:1 relationship between graph nodes and real-world intersections/dead-ends.
@@ 6. Conclusion
We argue that transport planners cannot expect accurate intersection counts from traditional methods and common data sources. Worse, these biases vary across cities and regions, thwarting consistent comparative urban analyses. To address this, our contribution is threefold: 1) topologically-corrected parsimonious models that better represent the real world; 2) more accurate street network measures—including intersection counts/densities, node degrees, and street segment lengths—for transport planning; 3) information compression to improve downstream graph algorithm memory use and runtime, improving analytical tractability. In an era of rapid urbanization, transport planners need good spatial models for evidence-informed planning. This article takes a step in that direction toward better models for better empirical science.
@@ Acknowledgments
The authors have no conflicts of interest to declare.

### v4

@@ 1. Introduction
Street intersections—the junctions where two or more street segments meet—are important transport network features whose configuration contributes to urban connectivity, centrality, compactness, walkability, and safety (Fogliaroni et al.,, 2018; Sharifi,, 2019; Cerin et al.,, 2022). Despite street intersections’ apparent simplicity, counting them can be nuanced and challenging in practice (Yang et al.,, 2020; Li et al.,, 2020). As we will argue, this is primarily due to three modeling challenges: network nonplanarity, intersection complexity, and curve digitization. In turn, intersection miscounts can bias downstream indicators of urban form, but this bias has not been fully accounted, no easily reusable algorithms exist to mitigate it, and intersection counts often remain ad hoc and naïve in research and practice.
This article measures this problem’s extent and asks: can we algorithmically simplify a street network graph’s topology to produce a more accurate and performant model? To answer this, it introduces and validates the OSMnx package’s reusable algorithms that ensure that graph nodes represent individual intersections or dead-ends and that graph edges represent individual street segments. This contribution offers three main advantages over existing techniques: (1) more accurate measures including intersection counts/densities, street segment lengths, node degrees, and betweenness centralities; (2) a topologically-corrected model that better represents the real world; and (3) information compression (i.e., fewer nodes/edges than in the original model) that can drastically speed-up subsequent network algorithms that scale with node or edge count.
@@ 2.1. Street Network Models
then we can say for $G$ that: (1) edge $(u,v)$ links from node $u$ to node $v$, (2) edge $(u,v)$ is incident to $u$ and to $v$, and (3) $v$ is a successor of $u$.
@@ 3.1. Edge Simplification Algorithm
OSMnx’s edge simplification algorithm solves the curve digitization problem, such as occurs with OpenStreetMap raw data as discussed earlier. It simplifies the graph’s topology by removing all nodes that are not “true” edge endpoints (i.e., intersections or dead-ends), then creating new (simplified) edges directly between those true endpoints and retaining the full geometries of the original edges as attributes on those new edges. The full algorithm can be found in the package (see endnotes), but we summarize it here.
@@
    If configured, either (a) the node possesses a user-defined attribute explicitly denoting it as an endpoint, or (b) the node’s incident edges possess a user-defined attribute and have different values than each other in that attribute.
@@ 3.3. Assessment and Validation
To validate the (topological) node consolidation algorithm, for each of these sampled urban areas we randomly sample 120 graph nodes, divided evenly between consolidated and not-consolidated nodes. Each node’s consolidation (or not) result is then validated by qualitatively cross-referencing it, by hand, against Google Street View and satellite imagery data. In comparing these resources, the validator asks whether the node(s) in question would be considered a single intersection by a pedestrian or motorist at that intersection. There are four categories of node consolidation validation result: (1) “true positive,” where the node should be consolidated and is, (2) “true negative,” where the node should not be consolidated and is not, (3) “false positive,” where the node should not be consolidated but is, and (4) “false negative,” where the node should be consolidated but is not. In keeping with the spirit of fuzzy sets, within the “true positive” category we create a subcategory for “true positive but insufficiently robust.” For example, if a sampled validation node is in a roundabout comprising five nodes that should have been consolidated into one, but the algorithm only consolidated four of those five, then this sampled node is assigned this result subcategory. In other words, this result improves on the original model and is mostly correct, but incomplete.
@@ 4. Validation and Assessment Results
Now the question becomes: are these reductions in node count legitimate, or merely false positives being incorrectly consolidated? To that end, the node consolidation algorithm’s validation can be considered from two perspectives, (1) the validity of the algorithm itself, and (2) the validity of its parameterization here. From the first perspective, this deterministic algorithm unsurprisingly performs with 100% accuracy—that is, each sampled node’s result aligns with expectations given this parameterization. From the second perspective, the algorithm performs imperfectly, but better than expected given this simplified constant parameterization per urban area used for global scalability in this preliminary demonstration.
Among nodes that were consolidated (i.e., the “positives”), the validation reveals a 98% true positive rate and a 2% false positive rate. Among these true positives, around 5% were “insufficiently robust,” meaning the algorithm was correct in consolidating a complex intersection but failed to fully merge all the nodes that the validator deemed integral during their qualitative assessment. Among nodes that were not consolidated (i.e., the “negatives”), the validation reveals an 87% true negative rate and a 13% false negative rate. In other words, our simplified parameterization here was likely overly conservative, leading to more false negatives than false positives through what is essentially algorithmic underfitting. This suggests that the extent of the intersection miscount problem could be even greater than our conservative assessment reveals.
@@ 5. Toward Better Models and Measures
Counting seems simple because its mathematics are simple. But in reality it can be difficult to define what to count and how to count it. Street network intersections exemplify this challenge. In transport planning, innumerable models, measures, and standards—from sustainable development certifications to resilience simulations—rely on intersection counts as input data. Accurate intersection densities, node degrees, and street segment lengths all depend on accurate street network models. Yet modeled intersections can be surprisingly difficult to count accurately, resulting in biased measures that propagate downstream. This study identified the sources of these biases, measured their prevalence, and presented and validated algorithms to better model street intersections for better data-driven, evidence-informed transport planning. These algorithms are generalizable but their implementation here is tied to one data source (OpenStreetMap) or others matching its structure. These open-source, freely reusable methods produce network models with a closer 1:1 relationship between graph nodes and real-world intersections/dead-ends.
@@ 6. Conclusion
We argue that transport planners cannot expect accurate intersection counts from traditional methods and common data sources. Worse, these biases vary across cities and regions, thwarting consistent comparative urban analyses. To address this, our contribution is threefold: (1) topologically-corrected parsimonious models that better represent the real world; (2) more accurate street network measures—including intersection counts/densities, node degrees, and street segment lengths—for transport planning; (3) information compression to improve downstream graph algorithm memory use and runtime, improving analytical tractability. In an era of rapid urbanization, transport planners need good spatial models for evidence-informed planning. This article takes a step in that direction toward better models for better empirical science.
@@ Acknowledgments
The author has no conflicts of interest to declare.